My employee seems to leave work anytime between 5PM and 6​PM, uniformly.

​a) What is the probability he will still be at work at 5:50PM?

​b) What is the probability he will still be at work at 5:50PM every day this week (M–​F)?

​c) What did you assume to calculate​ b?
Answer #1

Answer:

Given,

a = 5 PM

b = 6 PM

now consider,

difference = b - a

= 6 - 5

= 1 hour

= 60 minutes

x = 5:50 PM

x - a = 5:50 - 5 PM

= 50 minutes

a)

P(work at 5:50PM) = (x - a) / (b - a)

= 50/60

= 5/6

= 0.8333

b)

Required probability = (0.8333)^5

= 0.4018

c)

Here the probability of working time is independent from day of week.